# The Tianlai-WIYN North Celestial Cap Redshift Survey  [^1] Template AASArticle with Examples: v6.3.1Released on March, 1st, 2021

### v1

@@ V.1 SDSS extracted Mock Catalogs
SDSS mocks are created from the SDSS spectroscopic sample (2017AJ....154...28B) by selecting a set of a 3.5 radius disks on the sky centered on different directions, $(\alpha_{0},\delta_{0})$, on the sky. All objects within the chosen region, identified as galaxies (${\rm type}=3$), with a spectroscopic redshift $z<0.5$, and with an SDSS $g$ brighter than $g<19.5$, are rotated toward the NCP. A rotational transformation is determined for each disk which rotates the SDSS region to overlap the tNCCSz disk centered on the NCP, i.e. a transformation which rotates $(\alpha_{0},\delta_{0})$ to $\delta=90^{\circ}$. Each source is moved to the NCP region by this transformation. The $V$ band extinction ($A_{\mathrm{V}}$) at the rotated position is determined using the Planck extinction map. An un-extincted $V$ magnitude is computed from the SDSS $g$ magnitude using V=g0.3V=g-0.3. This 0.3-0.3 magnitude shift has been estimated using the color transformations of the SDSS photometric bands  [^6] See http://www.sdss3.org/dr8/algorithms/sdssUBVRITransform.php for SDSS to other photometric system conversion. (2006A&A...460..339J). The un-extincted $V$ is corrected for extinction by subtracting $A_{\mathrm{v}}$. A fraction of the sources are dropped, taking into account the NCCS selection efficiency (eq. 2) and the fiber assignment efficiency (eq. 3). This catalog of rotated $(\alpha,\delta)$’s and extincted $V$’s provides a mock catalog of tNCCSz targets. A subset of these targets are selected for the spectroscopic sample using the redshift determination efficiency function of eq. 4 to provide an tNCCSz mock catalog or “mock”. This procedure was repeated for different values of $(\alpha_{0},\delta_{0})$ to provide 16 tNCCSz mocks.
@@ V.2 Visual Comparison
Figure 11 also shows that within the small volume of the tNCCSz survey there is large variation in the numbers of galaxies, redshift distribution, and large scale structures. As with all pencil beam redshift surveys the dominant visual feature are dense clumps of galaxies along the line of sight separated by underdense regions 1991ApJ...379..482K. These are the intersection of the pencil beams with walls filaments and voids in the cosmic web.
@@ V.3.1 Random Catalogs
Estimators of $n$-point correlation functions (usually) make use of a “random catalog”. These are points given by a Poisson sampling of the survey volume after taking into account all selection functions and efficiencies. These points are meant to represent a distribution with no clustering, i.e. zero $n$-point correlation functions. We have generated such random tNCCSz galaxy catalogs using the efficiency functions defined in §IV.2. We start from the galaxy luminosity distribution defined as a Schechter function depending on the galaxy absolute $V$-band magnitude $M_{V}$ with the following parameters (see e.g 2001AJ....122..714B; 2009A&A...508.1217Z).
@@
Galaxies in the random catalog are initially uniformly distributed in the 3D space, with the galaxy number count proportional to the cosmological volume element, computed according to the standard $\Lambda CDM$ cosmology with Planck (2021A&A...652C...4P) parameters. The apparent NCCS magnitude is then computed, applying the distance modulus, and the Galactic extinction $Av$, without any $k$-correction. We then apply the NCCS survey selection efficiency (eq. 2), dropping a fraction of galaxies. We continue by applying the tNCCSz target selection cuts ($13<V_{\mathrm{mag}}<19$), followed by the declination dependent fiber assignment efficiency (equation 3). Finally, the galaxies for which we have a redshift estimates are selected applying the $\epsilon_{z}(V_{\mathrm{mag}})$ defined in equation 4.
@@ V.3.2 Landy-Szalay Estimators
Also plotted in Figure 12 is the statistical estimator mean and variance from the random catalog should be and is consistent with zero but only includes the statistical uncertainty from the finite number of galaxies. A much larger contribution to the variation in these estimators comes from sample variance of large scale structures in this relatively small and elongated cosmological volume. Large excursions from zero in the tNCCSz $\hat{\xi}_{\mathrm{LS}}$ are present at $d_{\mathrm{sep}}\sim 160,\,330\,\mathrm{Mpc}$. Such excursions are expected for pencil beam survey (see 1991ApJ...379..482K) because there are a small number of structures (voids, filaments etc.) of size comparable to the width of the survey along the line-of-sight. Different pencil beams have different structures leading to large excursions at different separations, as illustrated in this figure.
@@ V.5 Clusters
Despite the sparse sampling of our redshift catalog we have attempted to detect galaxy clusters using a simple Friends of Friends (FoF) cluster finder algorithm (1982ApJ...257..423H; 2016A&A...588A..14T). We allow a maximum redshift difference $\delta z=0.005$ ($1500\,\mathrm{km/s}$) and a transverse distance of $d_{\perp}=0.5\,\mathrm{Mpc}$ for cluster members. We require a minimum number of cluster members of $N^{\mathrm{gal}}_{\mathrm{min}}=3$ for cluster detection; this threshold increases to $N^{\mathrm{gal}}_{\mathrm{min}}=4$ for $0.03<z<0.06$ and $N^{\mathrm{gal}}_{\mathrm{min}}=5$ for $z<0.03$.
@@
To check the validity of our FoF cluster finder and quantify roughly our cluster sample purity, we detected clusters on 14 SDSS extracted mocks, half near $\delta=50^{\circ}$, and the other half centered at $\delta=25^{\circ}$. We have found an average of $\langle N_{\mathrm{clus}}\rangle\simeq 14$ per extracted mock, with a rather large dispersion as expected for pencil beam surveys. This number decreases to about $\langle N_{\mathrm{clus}}\rangle\simeq 8$ for $d_{\perp}=0.35\mathrm{Mpc}$. About 75% of all detected clusters have a redshift $z\lesssim 0.073$, corresponding to the volume limited published cluster catalog (2014yCat..35660001T), used as a reference or truth catalog. We cross matched each of our cluster catalog detected on SDSS mocks with the above mentioned reference catalog, and we have a matching fraction $\sim 85\%$ among the clusters in the redshift range covered by the reference catalog. The matching fraction drops to less than 10% if the extracted reference cluster catalogs are shuffled. As expected, the 2014yCat..35660001T catalog is richer than the list of clusters detected on SDSS extracted mocks and includes groups with only two galaxy members. Requiring minimal number of associated galaxies $(N_{\rm gal})$ as for the tNCCSz clusters, the average number of clusters found in the 2014yCat..35660001T catalog for each mock $\langle N_{\mathrm{clus}}^{\mathrm{Tempel}}\rangle\simeq 27$ is about twice the number of clusters identified by us in SDSS mocks.
@@ V.6 HI mass fraction
The motivation for this survey was to localize the HI in a survey volume of $\sim 3^{\circ}$ around the NCP out to $z\approx 0.07$. This volume will be surveyed for 21,cm emission by the Tianlai Dish Array described in 2021MNRAS.506.3455W. We have accomplished this localization in 3D with this optical redshift survey of galaxies, all of which will contain some HI. We have only determined the redshifts of a fraction of bright galaxies, those with $V<19$. Due to the high extinction near the NCP this corresponds to an extincted magnitude limit of $V\lesssim 18$. Such bright galaxies contain only a fraction of the HI. Thus, this survey only directly localizes a tracer of the total HI content of the survey volume. However, since the spatial distribution of this tracer will be strongly correlated that of the remainder of the HI containing galaxies we have, in a statistical sense, also indirectly localized all of the HI in the survey volume.
@@
We can estimate the fraction of HI contained in the sample of galaxies for which we have obtained redshifts (tNCCSz) by estimating the HI mass in each of these galaxies. Although we have no direct measurement of the HI content in any of our galaxies, we can use the ALFALFA survey Haynes:2011hi to find similar galaxies for which the HI mass has been measured. Specifically, for each of the tNCCSz galaxies ($Q_{\rm auto}\geq 2$) we correct the apparent VRI magnitudes for extinction and use the measured redshift ($z_{\rm auto}$) to determine their absolute VRI magnitudes. We then find the ALFALFA survey galaxy which is closest in absolute VRI space and assign the mass of that ALFALFA galaxy to the tNCCSz galaxy. This procedure gives a rough estimate of the HI mass that might be expected in the tNCCSz galaxy. We compare the estimated HI masses of the galaxies for which we have determined redshifts with the mean HI density of the universe at low redshifts. This is estimated by 10.1093/mnras/stz2038 to be $\Omega_{\mathrm{HI}}=0.0004$ or $\bar{\rho}_{\mathrm{HI}}=4.5\times 10^{7}\,M_{\odot}/\mathrm{Mpc}^{3}$. In Figure 15 we plot the cumulative HI mass in our survey volume as a function of redshift using 1) the estimated masses of the tNCCSz galaxies with measured redshifts, and 2) a uniform $\bar{\rho}_{\mathrm{HI}}$ density. The ratio of the two is a rough estimate of the fraction of HI contained in these galaxies. This ratio varies from $\sim 0.5$ at the lowest $z$ to $\sim 0.01$ at $z\sim 0.07$. This HI incompleteness is a function of both the galaxy redshift incompleteness at the magnitude limit and the magnitude limit itself - a deeper, more complete survey is better. Nevertheless, the HI which we have localized provides a good template for all the HI in the survey volume. This will be used to compare with future measurements by the Tianlai Dish Array survey described in 2022MNRAS.517.4637P.
@@
Figure 15: In this figure we plot the cumulative neutral hydrogen (HI) mass with redshift, i.e. the mass of HI with redshift less than $z$ in the survey area on the sky, $\delta\leq 87^{\circ}$. The green curve gives this cumulative distribution if all of the HI were spread out uniformly (see 10.1093/mnras/stz2038). The blue dots are rough estimates for the HI in the galaxies of the tNCCSz survey, with each dot representing the addition of a tNCCSz galaxy. The raggedness of this curve is real, representing the large inhomogeneities as a function of $z$ expected for pencil beam surveys (see 1991ApJ...379..482K). The same inhomogeneities should be present in the full distribution of HI including HI in galaxies not included in tNCCSz.
@@ VI Summary
In this paper we have presented the results of a redshift survey of galaxies with $m_{V}<19$ (including extinction) within $\sim 3^{\circ}$ of the North Celestial Pole (NCP) based on a published photometric catalog (Gorbikov_2014). Because of the proximity to the Galactic Plane, high extinction and difficulties associated with observations very near to the celestial poles no redshift survey of this region had been attempted until now. The motivation for this survey is to compare a future 21 cm (radio) redshift survey by the Tianlai project with an optical redshift survey. The non-tracking transit radio telescope can obtain the deepest exposure in a given amount of observation time at the NCP. Hence the NCP has advantages for radio observations even though it is not ideal for extra-Galactic optical astronomy.
@@
Several surveys observe the NCP region for the same reason the Tianlai project does – to obtain long, continuous observations of a limited region of the sky. Other HIM programs observe the NCP region, but to our knowledge only the Tianlai arrays can currently be tuned to overlap the redshift range of the galaxy redshift survey described here. However, this survey may also be useful for finding optical counterparts for transient radio phenomena, such as Fast Radio Bursts (FRBs). Several such radio transient surveys observe the NCP: the LOFAR Multifrequency Snapshot Sky Survey (MSSS) (2016MNRAS.456.2321S), the CHIME/FRB project (2021ApJS..257...59C).
@@ VII Acknowledgments
We thank Eric Hooper, Susan Ridgway, Heidi Schweiker, Dan Li, and Daryl Wilmarth for advice on using the WIYN telescope and Hydra spectrometer. This document was prepared by the Tianlai Collaboration and includes personnel and uses resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, HEP User Facility. Fermilab is managed by Fermi Research Alliance, LLC (FRA), acting under Contract No. DE-AC02-07CH11359. Work at UW-Madison and Fermilab was partially supported by NSF Award AST-1616554. Work at UW-Madison was further supported by an NSF REU award, the Graduate School, and the Thomas G. Rosenmeyer Cosmology Fund. Undergraduate students at Brown were supported by the Brown University SPRINT/UTRA program. Authors affiliated with French institutions acknowledge partial support from CNRS (IN2P3 & INSU), Observatoire de Paris and Irfu/CEA. They acknowledge financial support from “Programme National de Cosmologie and Galaxies” (PNCG) and from the FCPPL (France-China Particle Physics Laboratory) of CNRS, France. The Tianlai Dish Array was built with the support of the CAS special fund for repair and purchase, and operated with the support of the NAOC Astronomical Technology Center. The Tianlai Pathfinders are supported by the National SKA Program of China (Nos. 2022SKA0110100 and 2022SKA0110101), the National Natural Science Foundation of China (12361141814, 12203061, 12273070, 12303004).

### v2

@@ V.1 SDSS extracted Mock Catalogs
SDSS mocks are created from the SDSS spectroscopic sample (Blanton et al., 2017) by selecting a set of a 3.5 radius disks on the sky centered on different directions, $(\alpha_{0},\delta_{0})$, on the sky. All objects within the chosen region, identified as galaxies (${\rm type}=3$), with a spectroscopic redshift $z<0.5$, and with an SDSS $g$ brighter than $g<19.5$, are rotated toward the NCP. A rotational transformation is determined for each disk which rotates the SDSS region to overlap the tNCCSz disk centered on the NCP, i.e. a transformation which rotates $(\alpha_{0},\delta_{0})$ to $\delta=90^{\circ}$. Each source is moved to the NCP region by this transformation. The $V$ band extinction ($A_{\mathrm{V}}$) at the rotated position is determined using the Planck extinction map. An un-extincted $V$ magnitude is computed from the SDSS $g$ magnitude using V=g0.3V=g-0.3. This 0.3-0.3 magnitude shift has been estimated using the color transformations of the SDSS photometric bands  [^6] See http://www.sdss3.org/dr8/algorithms/sdssUBVRITransform.php for SDSS to other photometric system conversion. (Jordi et al., 2006). The un-extincted $V$ is corrected for extinction by subtracting $A_{\mathrm{v}}$. A fraction of the sources are dropped, taking into account the NCCS selection efficiency (eq. 2) and the fiber assignment efficiency (eq. 3). This catalog of rotated $(\alpha,\delta)$’s and extincted $V$’s provides a mock catalog of tNCCSz targets. A subset of these targets are selected for the spectroscopic sample using the redshift determination efficiency function of eq. 4 to provide an tNCCSz mock catalog or “mock”. This procedure was repeated for different values of $(\alpha_{0},\delta_{0})$ to provide 16 tNCCSz mocks.
@@ V.2 Visual Comparison
Figure 11 also shows that within the small volume of the tNCCSz survey there is large variation in the numbers of galaxies, redshift distribution, and large scale structures. As with all pencil beam redshift surveys the dominant visual feature are dense clumps of galaxies along the line of sight separated by underdense regions Kaiser and Peacock (1991). These are the intersection of the pencil beams with walls filaments and voids in the cosmic web.
@@ V.3.1 Random Catalogs
Estimators of $n$-point correlation functions (usually) make use of a “random catalog”. These are points given by a Poisson sampling of the survey volume after taking into account all selection functions and efficiencies. These points are meant to represent a distribution with no clustering, i.e. zero $n$-point correlation functions. We have generated such random tNCCSz galaxy catalogs using the efficiency functions defined in §IV.2. We start from the galaxy luminosity distribution defined as a Schechter function depending on the galaxy absolute $V$-band magnitude $M_{V}$ with the following parameters (see e.g Brown et al. (2001); Zucca et al. (2009)).
@@
Galaxies in the random catalog are initially uniformly distributed in the 3D space, with the galaxy number count proportional to the cosmological volume element, computed according to the standard $\Lambda CDM$ cosmology with Planck (Planck Collaboration et al., 2021) parameters. The apparent NCCS magnitude is then computed, applying the distance modulus, and the Galactic extinction $Av$, without any $k$-correction. We then apply the NCCS survey selection efficiency (eq. 2), dropping a fraction of galaxies. We continue by applying the tNCCSz target selection cuts ($13<V_{\mathrm{mag}}<19$), followed by the declination dependent fiber assignment efficiency (equation 3). Finally, the galaxies for which we have a redshift estimates are selected applying the $\epsilon_{z}(V_{\mathrm{mag}})$ defined in equation 4.
@@ V.3.2 Landy-Szalay Estimators
Also plotted in Figure 12 is the statistical estimator mean and variance from the random catalog should be and is consistent with zero but only includes the statistical uncertainty from the finite number of galaxies. A much larger contribution to the variation in these estimators comes from sample variance of large scale structures in this relatively small and elongated cosmological volume. Large excursions from zero in the tNCCSz $\hat{\xi}_{\mathrm{LS}}$ are present at $d_{\mathrm{sep}}\sim 160,\,330\,\mathrm{Mpc}$. Such excursions are expected for pencil beam survey (see Kaiser and Peacock (1991)) because there are a small number of structures (voids, filaments etc.) of size comparable to the width of the survey along the line-of-sight. Different pencil beams have different structures leading to large excursions at different separations, as illustrated in this figure.
@@ V.5 Clusters
Despite the sparse sampling of our redshift catalog we have attempted to detect galaxy clusters using a simple Friends of Friends (FoF) cluster finder algorithm (Huchra and Geller (1982); Tempel et al. (2016)). We allow a maximum redshift difference $\delta z=0.005$ ($1500\,\mathrm{km/s}$) and a transverse distance of $d_{\perp}=0.5\,\mathrm{Mpc}$ for cluster members. We require a minimum number of cluster members of $N^{\mathrm{gal}}_{\mathrm{min}}=3$ for cluster detection; this threshold increases to $N^{\mathrm{gal}}_{\mathrm{min}}=4$ for $0.03<z<0.06$ and $N^{\mathrm{gal}}_{\mathrm{min}}=5$ for $z<0.03$.
@@
To check the validity of our FoF cluster finder and quantify roughly our cluster sample purity, we detected clusters on 14 SDSS extracted mocks, half near $\delta=50^{\circ}$, and the other half centered at $\delta=25^{\circ}$. We have found an average of $\langle N_{\mathrm{clus}}\rangle\simeq 14$ per extracted mock, with a rather large dispersion as expected for pencil beam surveys. This number decreases to about $\langle N_{\mathrm{clus}}\rangle\simeq 8$ for $d_{\perp}=0.35\mathrm{Mpc}$. About 75% of all detected clusters have a redshift $z\lesssim 0.073$, corresponding to the volume limited published cluster catalog (Tempel et al., 2014), used as a reference or truth catalog. We cross matched each of our cluster catalog detected on SDSS mocks with the above mentioned reference catalog, and we have a matching fraction $\sim 85\%$ among the clusters in the redshift range covered by the reference catalog. The matching fraction drops to less than 10% if the extracted reference cluster catalogs are shuffled. As expected, the Tempel et al. (2014) catalog is richer than the list of clusters detected on SDSS extracted mocks and includes groups with only two galaxy members. Requiring minimal number of associated galaxies $(N_{\rm gal})$ as for the tNCCSz clusters, the average number of clusters found in the Tempel et al. (2014) catalog for each mock $\langle N_{\mathrm{clus}}^{\mathrm{Tempel}}\rangle\simeq 27$ is about twice the number of clusters identified by us in SDSS mocks.
@@ V.6 HI mass fraction
The motivation for this survey was to localize the HI in a survey volume of $\sim 3^{\circ}$ around the NCP out to $z\approx 0.07$. This volume will be surveyed for 21,cm emission by the Tianlai Dish Array described in Wu et al. (2021). We have accomplished this localization in 3D with this optical redshift survey of galaxies, all of which will contain some HI. We have only determined the redshifts of a fraction of bright galaxies, those with $V<19$. Due to the high extinction near the NCP this corresponds to an extincted magnitude limit of $V\lesssim 18$. Such bright galaxies contain only a fraction of the HI. Thus, this survey only directly localizes a tracer of the total HI content of the survey volume. However, since the spatial distribution of this tracer will be strongly correlated that of the remainder of the HI containing galaxies we have, in a statistical sense, also indirectly localized all of the HI in the survey volume.
@@
We can estimate the fraction of HI contained in the sample of galaxies for which we have obtained redshifts (tNCCSz) by estimating the HI mass in each of these galaxies. Although we have no direct measurement of the HI content in any of our galaxies, we can use the ALFALFA survey Haynes and others (2011) to find similar galaxies for which the HI mass has been measured. Specifically, for each of the tNCCSz galaxies ($Q_{\rm auto}\geq 2$) we correct the apparent VRI magnitudes for extinction and use the measured redshift ($z_{\rm auto}$) to determine their absolute VRI magnitudes. We then find the ALFALFA survey galaxy which is closest in absolute VRI space and assign the mass of that ALFALFA galaxy to the tNCCSz galaxy. This procedure gives a rough estimate of the HI mass that might be expected in the tNCCSz galaxy. We compare the estimated HI masses of the galaxies for which we have determined redshifts with the mean HI density of the universe at low redshifts. This is estimated by Hu et al. (2019) to be $\Omega_{\mathrm{HI}}=0.0004$ or $\bar{\rho}_{\mathrm{HI}}=4.5\times 10^{7}\,M_{\odot}/\mathrm{Mpc}^{3}$. In Figure 15 we plot the cumulative HI mass in our survey volume as a function of redshift using 1) the estimated masses of the tNCCSz galaxies with measured redshifts, and 2) a uniform $\bar{\rho}_{\mathrm{HI}}$ density. The ratio of the two is a rough estimate of the fraction of HI contained in these galaxies. This ratio varies from $\sim 0.5$ at the lowest $z$ to $\sim 0.01$ at $z\sim 0.07$. This HI incompleteness is a function of both the galaxy redshift incompleteness at the magnitude limit and the magnitude limit itself - a deeper, more complete survey is better. Nevertheless, the HI which we have localized provides a good template for all the HI in the survey volume. This will be used to compare with future measurements by the Tianlai Dish Array survey described in Perdereau et al. (2022).
@@
Figure 15: In this figure we plot the cumulative neutral hydrogen (HI) mass with redshift, i.e. the mass of HI with redshift less than $z$ in the survey area on the sky, $\delta\leq 87^{\circ}$. The green curve gives this cumulative distribution if all of the HI were spread out uniformly (see Hu et al. (2019)). The blue dots are rough estimates for the HI in the galaxies of the tNCCSz survey, with each dot representing the addition of a tNCCSz galaxy. The raggedness of this curve is real, representing the large inhomogeneities as a function of $z$ expected for pencil beam surveys (see Kaiser and Peacock (1991)). The same inhomogeneities should be present in the full distribution of HI including HI in galaxies not included in tNCCSz.
@@ VI Summary
In this paper we have presented the results of a redshift survey of galaxies with $m_{V}<19$ (including extinction) within $\sim 3^{\circ}$ of the North Celestial Pole (NCP) based on a published photometric catalog (Gorbikov and Brosch, 2014). Because of the proximity to the Galactic Plane, high extinction and difficulties associated with observations very near to the celestial poles no redshift survey of this region had been attempted until now. The motivation for this survey is to compare a future 21 cm (radio) redshift survey by the Tianlai project with an optical redshift survey. The non-tracking transit radio telescope can obtain the deepest exposure in a given amount of observation time at the NCP. Hence the NCP has advantages for radio observations even though it is not ideal for extra-Galactic optical astronomy.
@@
Several surveys observe the NCP region for the same reason the Tianlai project does – to obtain long, continuous observations of a limited region of the sky. Other HIM programs observe the NCP region, but to our knowledge only the Tianlai arrays can currently be tuned to overlap the redshift range of the galaxy redshift survey described here. However, this survey may also be useful for finding optical counterparts for transient radio phenomena, such as Fast Radio Bursts (FRBs). Several such radio transient surveys observe the NCP: the LOFAR Multifrequency Snapshot Sky Survey (MSSS) (Stewart et al., 2016), the CHIME/FRB project (CHIME/FRB Collaboration et al., 2021).
@@ VII Acknowledgments
We thank Eric Hooper, Susan Ridgway, Heidi Schweiker, Dan Li, and Daryl Wilmarth for advice on using the WIYN telescope and Hydra spectrometer. This document was prepared by the Tianlai Collaboration and includes personnel and uses resources of the Fermi National Accelerator Laboratory (Fermilab). This work was produced by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. Publisher acknowledges the U.S. Government license to provide public access under the DOE Public Access Plan DOE Public Access Plan. Work at UW-Madison and Fermilab was partially supported by NSF Award AST-1616554. Work at UW-Madison was further supported by an NSF REU award, the Graduate School, and the Thomas G. Rosenmeyer Cosmology Fund. Undergraduate students at Brown were supported by the Brown University SPRINT/UTRA program. Authors affiliated with French institutions acknowledge partial support from CNRS (IN2P3 & INSU), Observatoire de Paris and Irfu/CEA. They acknowledge financial support from “Programme National de Cosmologie and Galaxies” (PNCG) and from the FCPPL (France-China Particle Physics Laboratory) of CNRS, France. The Tianlai Dish Array was built with the support of the CAS special fund for repair and purchase, and operated with the support of the NAOC Astronomical Technology Center. The Tianlai Pathfinders are supported by the National SKA Program of China (Nos. 2022SKA0110100 and 2022SKA0110101), the National Natural Science Foundation of China (12361141814, 12203061, 12273070, 12303004).